In 2003, Adam was 24 years old. In 2006 Adam will be three times as old as eve. How old was eve in 2003?
Answer by addingup(3677)

So, in 2006 Adam is 2006-2003=3 years older than 24; 24+3= 27. And at that point, he is 3 times older than Eve. Let's call Eve E:
24+3= 3E <--Adam in 2006 is 24+3= 27 years old. And 3 x Eve= 27.
27= 3E
E= 9 <-- this is in 2006, Eve is 9 years old
9-3= 6 <-- This is Eve in 2003, she's 9-3= 6 years old.